Spray drift is a constant concern for pesticide applicators and agricultural producers. The inverse relationship between droplet size and drift potential is well known. A researcher is interested in the effects of herbicide formulation on spray atomization. The researcher believes a normal distribution with mean 1050 µm and standard deviation 150 µm is a reasonable model for droplet size for water (the "control treatment") sprayed through a 760 ml/min nozzle. (a) What is the probability that the size of a single droplet is less than 1485 µm? Greater than 1000 µm? (Round your answers to four decimal places.) less than 1485 µm greater than 1000 µm (b) What is the probability that the size of a single droplet is between 1000 and 1485 µm? (Round your answer to four decimal places.) (c) How would you characterize the smallest 2% of all droplets? (Round your answer to two decimal places.) The smallest 2% of droplets are those smaller than µm in size. (d) If the sizes of five independently selected droplets are measured, what is the probability that at least one exceeds 1485 µm? Hint: See Binomial Distribution for Lesson 6. (Round your answer to four decimal places.)
Answer #1

a) The probability that the size of a single droplet is less than 1485 µm is

P(X < 1485) = P(Z < (1485-1050)/150) = P(Z < 2.9) = 0.99813

The probability that the size of a single droplet is greater than 1000 µm is

P(X > 1000) = P(Z > (1000-1050)/150) = P(Z >-0.3333) = 0.63056

b) The probability that the size of a single droplet is between 1000 and 1485 µm is

P(1000 <X< 1485) = P(  (1000-1050)/150 < Z < (1485-1050)/150) = P(-0.3333 < Z < 2.9) = P(Z < 2.9) - P(Z <-0.3333) = 0.99813 - 0.36945 = 0.62868

c) The smallest 2% of all droplets is X = mean - Zalpha SD = 1050 -2.054 *150 = 741.9

d) Given n = 5, The probability that the size of a single droplet is less than 1485 µm is 0.99813

The probability that the size of a single droplet is greater than 1485 µm is 1 - 0.99813 = 0.00187

The probability that at least one exceeds 1485 µm is

P(X 1) = 1-P(X < 1) = 1-Ί .|(0.00187) (1-000187)5-0 = 1-0.990685= 0.009315

> B) & A) incorrect, because z > -0.333 from the Z- table it is = 0.3707

Akaram Tue, Mar 29, 2022 9:08 PM